# FovealNet: Advancing AI-Driven Gaze Tracking Solutions for Optimized Foveated Rendering System Performance in Virtual Reality

### v1

@@ Introduction
  • To reduce the implementation cost of FovealNet, we introduce an efficient input cropping method that focuses on extracting the central eye region, allowing for effective reuse of gaze-tracking results. Furthermore, we propose an adaptive input token pruning technique for the transformer-based gaze-tracking DNN, achieving superior computational efficiency while maintaining the tracking performance.
@@ 3 FovealNet Design
The overall structure of FovealNet is illustrated in Fig. 6. When an eye image is received, it is initially compared with previous buffer frames to determine if the earlier gaze prediction results can be reused, as detailed in Section 3.1. If the previous results are not applicable, the eye image is cropped using the method outlined in Section 3.1. The cropped image is then passed to the gaze tracking ViT, which is explained in Section 3.2 and Section 3.3. Finally, the multi-resolution training approach is described in Section 3.4.
@@ 3.1 Event-based Cropping for Efficient TFR
Eye images captured by near-eye cameras often contain redundant pixels (e.g., background, facial muscles) that are irrelevant for gaze tracking prediction. These pixels can negatively impact the prediction results and increase the computational cost. To tackle this, we propose an event-based analytical approach that efficiently crops the informative regions of the eye. Additionally, since eye frames are streamed into the TFR system consecutively and share content similarities, we can reuse the gaze prediction results of previous frames, further reducing computational cost.
Input: Current and previous frames $F^{t}$, $F^{t-1}$, buffered gaze $G$, activation threshold $\beta_{1}$, reuse threshold $\beta_{2}$, Kernel $B$, pupil detect function $is\_pupil$.
Output: Gaze direction $G$
1
2 Initiation
3       $M_{ij}\leftarrow\sigma\left(\frac{\Delta F_{ij}^{t}}{F_{ij}^{t-1}},\beta_{1}\right)$;
4       if $\sum_{ij}M_{i,j}<\beta_{2}$ then
             return $G$;
              // Reuse previous gaze
5
6      else
7             $C\leftarrow invert\_binarize(F^{t})$ ;
             $C\leftarrow(C\ominus B)\oplus B$ ; // Find the LCC within C.
             // Check the Shape of LCC
8             if $is\_pupil(C)$ then
9                   Cropping and gaze tracking ;
10                   $G\leftarrow\text{Predicted Gaze}$ ;
11                   $F^{t-1}\leftarrow F^{t}$;
                   return $G$;
                    // Return new gaze direction
12
13            else
                   return $G$;
                    // Reuse previous direction
14
Algorithm 1 FovealNet Overall Workflow
@@ 3.1.1 Region Cropping Algorithm
3.1.2 Gaze Reuse Criterion
Usually the eye movements is small, resulting in no significant change in gaze direction [60, 68]. Therefore we can reuse the previous gaze direction result for save computation. To achieve this, we use the activated pixels in the event map as a basis for the decision.
The event map is obtained by processing consecutive near-eye image inputs. First, we calculate the pixel-wise inter-frame difference $\Delta F^{t}$ between two successive frames $F^{t-1}$ and $F^{t}$. Next, each spatial element $\Delta F_{ij}^{t}\in\Delta F_{t}$ is normalized using the corresponding values in $F_{ij}^{t-1}$. Based on the normalized values, we determine whether each pixel should be represented in the event map. Finally, we compare the number of activated pixels with a predefined threshold to decide whether the current frame can reuse the previous gaze direction result. This process can be formulated as:
$\sigma\left(\sum_{i,j}\sigma\left(\frac{\Delta F_{ij}^{t}}{F_{ij}^{t-1}},\beta_{1}\right),\beta_{2}\right)$ (1)
where $\sigma(x,\beta)$ denotes a step function that returns 1 if $x>\beta$ and 0 otherwise. The overall activation of the event map is evaluated by other step function which uses $\beta_{2}$ as its threshold. If the total number of activated pixels exceeds $\beta_{2}$, gaze reuse is considered unfeasible; otherwise, it is deemed feasible. To mitigate cumulative errors, if the gaze direction from the previous frame is reused, the frame $F_{t-1}$ remains unchanged. Conversely, $F_{t-1}$ is updated to $F_{t}$. Both $\beta_{1}$ and $\beta_{2}$ are hyper-parameters that can be fine-tuned for specific applications and scenarios.
Algorithm 1 provides a detailed explanation of the region cropping and gaze reuse criteria. Unlike previous method [11] that uses a DNN to predict the cropped region. Our approach is purely analytical based whose computation cost is only 21.6% of [11] (11.98M FLOPs vs. 55.4M FLOPs). Additionally, our approach is particularly effective in low-motion scenarios, enabling efficient reuse of previously detected gaze results and reducing the need to activate the tracking DNN for every frame. This leads to improved performance with minimal computational cost.
@@ 3.2 Efficient Gaze Tracking Network
Figure 8: The design of the gaze tracking ViT with tokenwise pruning mechanism.
The cropped eye images containing informative content are first resized to a smaller square (224×224) and then processed by the gaze tracking DNN to predict gaze direction. FovealNet employs a vision transformer (ViT) architecture [30] for this task, as it provides superior performance compared to convolutional neural network (CNN)-based architectures. The ViT operates by dividing the input image into patches, which are then tokenized and appended with positional information, the outputs are then passed through the transformer block. The ViT contains 8 transformer block, each block consists of 6 heads with an embedding dimension of 384. We also modify the original ViT by replacing the classifier MLP layers with a sequence of linear layers to output the 2D gaze direction. A completed version of gaze tracking ViT is depicted in Fig. 8.
A key advantage of ViT over CNN is its ability to fine-grain prune input tokens, enabling the removal of image tokens with unimportant content, as shown in Fig. 2. In the self-attention mechanism, tokens are linearly transformed into Query, Key, and Value matrices. The attention score is then computed by performing a dot product between the Query and Key matrices, followed by scaling and Softmax operation. The attention score reflects the importance of each token in relation to the gaze prediction result. Using these scores, we employ a top-k selector to remove unimportant tokens, which further reduces the computational cost of subsequent ViT blocks by decreasing the number of input tokens, outlined in Fig. 8.
@@ 4.1 Settings
In the data pre-processing stage, we first horizontally flip the right eye images to align them with the left eye images, a common practice that allows the application of a single mapping for both eyes [23]. We apply $\beta_{1}=0.2$ and $\beta_{2}=500$ for cropping the input image in Eq. 1. To train the gaze tracking ViT, we use a series of data augmentations techniques to enhance the training convergence speed. These techniques include random cropping, where the image is cropped to a random size between $80\%$ to $100\%$ of its original size, followed by a random shift in position by up to $10\%$ of the image’s width and height. Finally, we normalize the images to standardize the input data for the model.

### v2

@@ Introduction
  • To reduce the implementation cost of FovealNet, we introduce an efficient input cropping method that focuses on extracting the central eye region. Furthermore, we propose an adaptive input token pruning technique for the transformer-based gaze-tracking DNN, achieving superior computational efficiency while maintaining the tracking performance.
@@ 3 FovealNet Design
The overall structure of FovealNet is illustrated in Fig. 6. When an eye image is received, the eye image is cropped using the method outlined in Section 3.1. The cropped image is then passed to the gaze tracking ViT, which is explained in Section 3.2 and Section 3.3. Finally, the multi-resolution training approach is described in Section 3.4.
@@ 3.1 Event-based Cropping for Efficient TFR
Eye images captured by near-eye cameras often contain redundant pixels (e.g., background, facial muscles) that are irrelevant for gaze tracking prediction. These pixels can negatively impact the prediction results and increase the computational cost. To tackle this, we propose an event-based analytical approach that efficiently crops the informative regions of the eye.
@@ 3.2 Efficient Gaze Tracking Network
The cropped eye images containing informative content are first resized to a smaller square (224×224) and then processed by the gaze tracking DNN to predict gaze direction. FovealNet employs a vision transformer (ViT) architecture [30] for this task, as it provides superior performance compared to convolutional neural network (CNN)-based architectures. The ViT operates by dividing the input image into patches, which are then tokenized and appended with positional information, the outputs are then passed through the transformer block. The ViT contains 8 transformer block, each block consists of 6 heads with an embedding dimension of 384. We also modify the original ViT by replacing the classifier MLP layers with a sequence of linear layers to output the 2D gaze direction.
A key advantage of ViT over CNN is its ability to fine-grain prune input tokens, enabling the removal of image tokens with unimportant content, as shown in Fig. 2. In the self-attention mechanism, tokens are linearly transformed into Query, Key, and Value matrices. The attention score is then computed by performing a dot product between the Query and Key matrices, followed by scaling and Softmax operation. The attention score reflects the importance of each token in relation to the gaze prediction result.
@@ 4.1 Settings
In the data pre-processing stage, we first horizontally flip the right eye images to align them with the left eye images, a common practice that allows the application of a single mapping for both eyes [23]. We apply $\beta_{1}=0.2$ and $\beta_{2}=500$ for cropping the input image in LABEL:eqn:cropping. To train the gaze tracking ViT, we use a series of data augmentations techniques to enhance the training convergence speed. These techniques include random cropping, where the image is cropped to a random size between $80\%$ to $100\%$ of its original size, followed by a random shift in position by up to $10\%$ of the image’s width and height. Finally, we normalize the images to standardize the input data for the model.